OpenStudy (anonymous):

Locate the absolute extrema of the function of the closed interval. y= (4/x) + (tan((pii*x)/8)), [1,2]

OpenStudy (anonymous):

\[y = \frac{4}{x}+\tan(\frac{\pi x}{8}); x\in [1,2]\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\frac{dy}{dx} = -\frac{4}{x^2} + (\sec^2(\frac{\pi x}{8})*\frac{\pi}{8})=0\]

OpenStudy (anonymous):

ohh! i almost had it. i just didn't know what to do with part of the tangent function

OpenStudy (anonymous):

hmm there doesn't seem to be any absolute extrema in [1,2] :-(